OpenStudy (anonymous):

Find the simple interest rate on a $400 investment that grows to $425 in 9 months. Rate____________________

OpenStudy (anonymous):

I = Prt 425-400 = 400 r (9/12) r = 25 / (400 x .75) r = 0.0833333333 or r = 8.3333 %